### v1

@@ V Existence of the Minimum
The first term goes to 0 since $PM_{f}$ is a compact operator. The second term is contained in $C\epsilon,C\epsilon]$ for $n$ large enough where the constant $C$ depends only on $f$. Since $\epsilon$ can be arbitrarily small we conclude that $\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})M_{f}\}\to 0$ and hence $\Gamma_{0}\to\rho$. By Eq. (23) we get $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{0}(H_{0}+\beta^{-1}\log\Gamma_{0})\}$ which concludes the proof. ∎
@@
If we combine Theorem 7 and Theorem 11, we get that each density $\rho\in\mathscr{X}_{>0}$ is $v$-representable, which is the first part of Theorem 1. The full proof for Theorem 1 can be found in Section VIII.
@@ VI Closing the Circle
So far we showed that each density in $\mathscr{X}_{>0}$ is $v$-representable. In this section we show that each density coming from a Gibbs state is contained in $\mathscr{X}_{>0}$. At elevated temperatures, all excited state of the Hamiltonian contribute to the Gibbs state (and hence also to the density) and since these state form an orthonormal basis, it is not possible that the density vanishes somewhere. We formulate this result in the following theorem that closes the circle from strictly positive densities in $\mathscr{X}_{>0}$ to representing potentials in $\mathscr{X}^{*}$ and back to a strictly positive density of the corresponding Gibbs state.
@@
The density of any Gibbs state is everywhere non-zero.
@@
where the $\lambda_{j}$ are the eigenvalues of $H_{v}$ with corresponding eigenstates $\psi_{j}$. We denote the density of $\Gamma_{v}$ by $\rho$ and the densities of $\psi_{j}$ by $\rho_{j}$. Let us assume that the density $\rho$ is equal to 0 at $x_{0}$. By the relation
@@
we conclude that $\rho_{j}(x_{0})=0$ for all $j$ since the weights are all non-zero. This implies that all $\psi_{j}$ vanish on $U=\{x_{0}\}\times[0,1]^{N-1}$. (The restriction of $\psi_{j}$ to $U$ is well-defined by the trace operator for Sobolev spaces [25, p. 315].) Consequently, each linear combination of the $\psi_{j}$ vanishes on $U$. Yet, this statement makes only sense if the linear combination is in $H^{1}$ too. Since the $\psi_{j}$ form a basis of $L^{2}(\mathbb{T}^{N})$ we also need to be able to express each function in $H^{1}(\mathbb{T}^{N})\subset L^{2}(\mathbb{T}^{N})$ as their linear combination. Now this yields a contradiction since clearly not every function in $H^{1}(\mathbb{T}^{N})$ vanishes on U. We note that while the restriction to $U$ is well-defined for $H^{1}(\mathbb{T}^{N})$ functions, this is not the case for functions in $L^{2}(\mathbb{T}^{N})$. ∎
@@ VII Gâteaux differentiability
The last part of the main result (Theorem 1) is about Gâteaux differentiability of $F_{\mathrm{DM}}^{\beta}$ at densities $\rho\in\mathscr{X}_{>0}$. As mentioned in Section IV, Gâteaux differentiability can be achieved by continuity and a single-element subdifferential. Thus, the goal of this section is to show that each density $\rho\in\mathscr{X}_{>0}$ has only one subgradient. We do that by using standard results from convex analysis and by exploiting the following lemma.
Lemma 13.
The Helmholtz free energy $\Omega^{\beta}(v)$ is a strictly concave function in $v\in H^{-1}(\mathbb{T})$.
This result is obtained by exploiting the uniqueness of the Gibbs state. The proof proceeds analogously to [16, Th. 10]. Additionally, we need the following definition and lemmata.
Definition 14 (Barbu and Precupanu [23], p. 71).
Let $f$ be a convex function on $\mathscr{X}$. The closure of $f$, denoted by $\mathrm{cl}f$, is the lower semi-continuous hull of $f$, i.e.  $\mathrm{cl}f(\rho)=\liminf_{\rho^{\prime}\to\rho}f(\rho^{\prime})$ for all $\rho\in\mathscr{X}$ if $\liminf_{\rho^{\prime}\to\tilde{\rho}}f(\rho^{\prime})>$-$\infty$ for all $\tilde{\rho}$ and it is the constant function -$\infty$ if there is a $\tilde{\rho}\in\mathscr{X}$ such that $\liminf_{\rho^{\prime}\to\tilde{\rho}}f(\rho^{\prime})=$-$\infty$.
Since the universal functional $F_{\mathrm{DM}}^{\beta}$ is uniformly bounded from below by Lemma 6, the closure of it is given by $\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})$. Another characterization of the closure of a convex function is by its biconjugate as the following lemma shows.
Lemma 15 (Barbu and Precupanu [23], Cor. 2.23).
The biconjugate of a convex function $f$ coincides with its closure. That is, $f^{**}=\mathrm{cl}f$.
Lemma 16 (Barbu and Precupanu [23], Prop. 2.33).
Let $f:\mathscr{X}\to\mathbb{R}\cup\{\infty\}$ be a proper, convex and lower semi-continuous function and $f^{*}$ its conjugate. Then the following two statements are equivalent,
  1. (i)
    $v\in\partial f(\rho)$
  2. (ii)
    $\rho\in\partial f^{*}(v)$.
@@
We have now all the needed ingredients to show Gâteaux differentiability of the universal functional.
@@
The universal functional $F_{\mathrm{DM}}^{\beta}$ is Gâteaux differentiable at any density $\rho\in\mathscr{X}_{>0}$.
@@
According to Theorem 7, the functional $F_{\mathrm{DM}}^{\beta}$ is continuous on $\mathscr{X}_{>0}$. By Lemma 17 it is then enough to show that $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is single-valued for all densities $\rho\in\mathscr{X}_{>0}$. The conjugate function of $F_{\mathrm{DM}}^{\beta}$ is given by $\Omega^{\beta}(-\cdot)$. If we conjugate this function we end up with the biconjugate of $F_{\mathrm{DM}}^{\beta}$ which is defined by
$F_{\mathrm{DM}}^{\beta**}(\rho)=\sup\big{\{}\Omega^{\beta}(v)-\langle\rho,v\rangle\mid v\in H^{-1}(\mathbb{T})\big{\}}.$ (32)
The biconjugate $F_{\mathrm{DM}}^{\beta**}$ is a convex, lower semi-continuous function and it coincides with the convex envelope of $F_{\mathrm{DM}}^{\beta}$. Since $F_{\mathrm{DM}}^{\beta}$ is convex, Lemma 15 states that the closure of $F_{\mathrm{DM}}^{\beta}$ equals to its biconjugate $F_{\mathrm{DM}}^{\beta**}$. By Definition 14 and its proceeding discussion $\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf\limits_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})$. Theorem 7 shows that the universal functional $F_{\mathrm{DM}}^{\beta}$ is continuous on $\mathscr{X}_{>0}$ and hence we get that for each $\rho>0$
$F_{\mathrm{DM}}^{\beta**}(\rho)=\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf\limits_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})=F_{\mathrm{DM}}^{\beta}(\rho).$ (33)
This means that $F_{\mathrm{DM}}^{\beta}$ and its biconjugate $F_{\mathrm{DM}}^{\beta**}$ coincide on the open set $\mathscr{X}_{>0}$. In particular this means that the subdifferentials are equal for densities $\rho\in\mathscr{X}_{>0}$, i.e. $\partial F_{\mathrm{DM}}^{\beta}(\rho)=\partial F_{\mathrm{DM}}^{\beta**}(\rho)$. So instead of showing that $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is single-valued, we show it for $\partial F_{\mathrm{DM}}^{\beta**}(\rho)$. Note that the conjugate function of $F_{\mathrm{DM}}^{\beta**}$ is $\Omega^{\beta}(-\cdot)$ since the triple conjugate of $F_{\mathrm{DM}}^{\beta}$ is equal to the conjugate. The function $F_{\mathrm{DM}}^{\beta**}$ fulfills all assumptions of Lemma 16 and hence we get
$v\in\partial F_{\mathrm{DM}}^{\beta**}(\rho)\Longleftrightarrow\rho\in\partial\left(-\Omega^{\beta}\right)(v).$ (34)
By Lemma 13 we have that $\Omega^{\beta}$ is strictly convex. In particular this means that a density $\rho$ can only be contained in one subdifferential. Otherwise, if $\rho$ is contained in the subdifferential at $v_{1}$ and $v_{2}$, then the function $\Omega^{\beta}$ would not be strictly convex on the line with boundary points $v_{1}$, $v_{2}$. By the relation (34) we get that $\partial F_{\mathrm{DM}}^{\beta**}(\rho)$ is single valued. This concludes the proof. ∎
@@ VIII Summary and outlook
Let $\rho\in\mathscr{X}_{>0}$, then Theorem 7 shows that $\partial F^{\beta}_{\mathrm{DM}}(\rho)$ is non-empty. This implies that $\rho$ is a minimizer for the variational principle in Eq. (15a), where the negative potential is in the subdifferential, i.e., $v\in\partial F^{\beta}_{\mathrm{DM}}(\rho)$. The minimizer of Eq. (15b), guaranteed to exist by Theorem 11, is then also a minimizer in the Helmholtz functional (Eq. (12)). This minimizer is unique by Proposition 3 and is equal to the Gibbs state $\Gamma_{v}$. We conclude, that $\rho$ is the density of $\Gamma_{v}$ and hence $v$-representable. On the other hand, every density of a Gibbs state $\Gamma_{v}$ with $v\in H^{-1}(\mathbb{T})$ is contained in $\mathscr{I}$ according to Proposition 4 and is strictly positive by Theorem 12. In particular, every $v$-representable density is in $\mathscr{X}_{>0}$ and $F^{\beta}_{\mathrm{DM}}$ is not differentiable at densities in $\mathscr{I}\setminus\mathscr{X}_{>0}$. On the other hand, Theorem 18 shows that the universal functional $F^{\beta}_{\mathrm{DM}}$ is Gâteaux differentiable at any $\rho\in\mathscr{X}_{>0}$. ∎
@@
In this work we provided an explicit characterization of the set of uniquely $v$-representable densities on the one-dimensional torus at elevated temperatures and established (Gâteaux) differentiability of the universal functional. In the same manner the theory can be developed for any other one-dimensional bounded domain. As an extensive parameter, the volume needs to be finite, and thus a periodic setting is the natural choice for elevated temperatures. The restriction to one dimension remains as the main limitation of the current formulation. Another interesting observation is that we do need distributional potentials. In DFT often the potential set $L^{3/2}+L^{\infty}$ is considered [6]. However, for zero temperature and one particle it was already observed by Englisch and Englisch [26] and Chayes et al. [12] that relatively nice densities need distributional potentials. Sutter et al. [7] showed that distributional potentials are sufficient for $v$-representability of the set $\mathscr{X}_{>0}$. For elevated temperature we further show that distributional potential are also necessary. This leads to the clear conjecture that also for higher dimensional domains some sort of distributional potentials are needed. However, the continuous Sobolev embedding $H^{1}\hookrightarrow L^{\infty}$ is crucial and it holds only for one-dimensional domains. For higher dimension we either need to require higher regularity or higher integrability conditions. This means that for the Sobolev space $W^{m,p}$, we need to have $m$ and $p$ large enough. The problem is then that not all Gibbs state densities might be contained in $W^{m,p}$ and that a Hamiltonian with a potential in $W^{-m,q}=(W^{m,p})^{\ast}$ might not be self-adjoint. Another open question is how to extend the results to the grand canonical ensemble of variable particle number. In this case, the particle-number operator is not bounded anymore and this is currently needed in the proof of Theorem 11 where we show that the weak-* limit yields the same density as the sequence. Additionally, we do not know if the square root of the Gibbs state density is in $H^{1}$. However, we think that the rest of the theory does not fundamentally change and that we get the same conclusion as for the canonical ensemble.
@@ Acknowledgements.
MP acknowledges support from the German Research Foundation under Grant SCHI 1476/1-1 and from ERC-2021-STG grant agreement No. 101041487 REGAL. The authors SMS and KJHG thank The Netherlands Organization for Scientific Research, NWO, for its financial support under Grant No. OCENW.KLEIN.434 and Vici Grant No. 724.017.001. RvL acknowledges the Academy of Finland grant under project number 356906. MR acknowledges the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG), EXC 2056, project ID 390715994.
@@ Appendix A Proof of Proposition 3
We still need to show that Proposition 3 holds. The result that the unique minimizer of the Helmholtz functional is given by the Gibbs state $\Gamma_{v}=\mathrm{e}^{\beta H_{v}}/Z(v)$ and the relation $\Omega^{\beta}_{v}(\Gamma_{v})=-\beta\log Z(v)$ is shown in standard literature. Yet, we give here a different approach than Mermin [11] and a more elaborated version than our previous works [16; 15].
@@
To determine the minimizer of the Helmholtz functional $\Omega^{\beta}_{v}(\Gamma)=\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$ we make it stationary with respect to small perturbations $\delta\Gamma$ that obey $\operatorname{Tr}\{\delta\Gamma\}=0$. We compute
@@
For the last remaining part of the proposition we need to show that $Z(v)$ is indeed finite for any potential $v\in H^{-1}(\mathbb{T})$. First, we show that the partition function is finite for a non-interacting system with zero potential, i.e., $H_{v}=T$. We denote this partition function by $Z_{0}(0)$. The eigenstates of the kinetic-energy operator $\frac{1}{2}\Delta$ on $\mathbb{T}$ are Slater determinants built from the functions $\mathrm{e}^{\mathrm{i}2\pi p\cdot x}$ with $p\in\mathbb{Z}$ which have eigenvalues $2\pi^{2}p^{2}$. We estimate
@@
To show that $Z(v)$ is finite for a general interaction and a general potential we first show that $\Omega^{\beta}(v)$ is finite. For that recall that we assume that the interaction satisfies the KLMN condition (3). Moreover, Sutter et al. [7, Lem. 16] showed that any potential $v\in H^{-1}$ fulfills also the KLMN condition $|V(\Psi,\Psi)|\leq aT(\Psi)+b$ where $a>0$ can be made arbitrary small. Here $V(\cdot,\cdot)$ is the quadratic form produced by $v\in H^{-1}$. Combing these two KLMN conditions guarantees the existence of $0<a<1$ and $b$ such that $\langle\Psi,H_{v}\Psi\rangle\geq(1-a)T(\Psi)-b$ for all $\Psi\in\mathscr{W}$. This condition translates to density matrices and we get
$\Omega^{\beta}(v)=\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log(\Gamma))\}\geq(1-a)\inf_{\Gamma}\operatorname{Tr}\{\Gamma(T+\big{(}\beta(1-a)\big{)}^{-1}\log(\Gamma))\}-b=-\beta^{-1}\log\big{(}Z_{0}(0)\big{)}-b>-\infty.$ (39)
Let $|H_{v}=\sum_{j=1}^{\infty}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of the Hamiltonian $H_{v}$ which exist due to the compact embedding of $Q(H_{v})=\big{(}H^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\big{)}^{\wedge N}$ into $\big{(}L^{2}(\mathbb{T})\otimes\mathbb{C}^{2}\big{)}^{\wedge N}$. We define a sequence of density matrices by
@@
The left-hand side is bounded from below by Eq. (39) and we get that $Z_{n}[v]$ is uniformly bounded from above. This implies $Z(v)<\infty$. ∎

### v2

@@ V Existence of the Minimum
The first term goes to 0 since $PM_{f}$ is a compact operator. The second term is contained in $C\epsilon,C\epsilon]$ for $n$ large enough where the constant $C$ depends only on $f$. Since $\epsilon$ can be arbitrarily small we conclude that $\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})M_{f}\}\to 0$ and hence $\Gamma_{0}\mapsto\rho$. By Eq. (30) we get $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{0}(H_{0}+\beta^{-1}\log\Gamma_{0})\}$ which concludes the proof. ∎
@@
If we combine Theorem 7 and Theorem 11, we get that $\rho\in\mathscr{X}_{>0}$ is a sufficient condition for $v$-representability by a Gibbs state, which is the first part of Theorem 1. The full proof for Theorem 1 can be found in Section VIII.
@@ VI Closing the Circle
So far we showed that each density in $\mathscr{X}_{>0}$ is $v$-representable by a Gibbs state. In this section we show that each $v$-representable density is strictly positive and thus contained in $\mathscr{X}_{>0}$. Consequently, the restriction to strictly positive densities in order to show $v$-representability of a density is necessary. In order to show this, observe that at elevated temperatures, all excited states of the Hamiltonian contribute to the Gibbs state (and hence also to the density) and since these states form an orthonormal basis, it is not possible that the density vanishes somewhere. We formulate this result in the following theorem that closes the circle from strictly positive densities in $\mathscr{X}_{>0}$ to representing potentials in $\mathscr{X}^{*}$ and back to a strictly positive density of the corresponding Gibbs state.
@@
The density of any Gibbs state is strictly positive.
@@
where the $\lambda_{j}$ are the eigenvalues of $H_{v}$ with corresponding eigenstates $\psi_{j}\in\mathscr{W}$. We denote the density of $\Gamma_{v}$ by $\rho$ and the densities of $\psi_{j}$ by $\rho_{j}$. Let us assume that the density $\rho$ is equal to 0 at $x_{0}$. By the relation
@@
we conclude that $\rho_{j}(x_{0})=0$ for all $j$ since the weights are all non-zero. This implies that all $\psi_{j}$ vanish on $U=\{x_{0}\}\times[0,1]^{N-1}\times\{\uparrow,\downarrow\}^{N}$. Note that the restriction to $U$ of $\psi_{j}\in\mathscr{W}$ or any linear combination of those vectors in $(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}$ is well-defined by the trace operator for Sobolev spaces [4, p. 315], which is not the case for a general element in $\mathcal{H}=(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}$. Since the $\psi_{j}$ form a basis of $(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}$ we also need to be able to express each function in $(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}\subset(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}$ as their linear combination. Now this yields a contradiction since clearly not every function in $(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}$ vanishes on $U$. ∎
@@ VII Gâteaux differentiability
The last part of the main result (Theorem 1) is about Gâteaux differentiability of $F_{\mathrm{DM}}^{\beta}$ at densities $\rho\in\mathscr{X}_{>0}$. As mentioned in Section IV, Gâteaux differentiability can be achieved by continuity and a single-valued subdifferential. Thus, the goal of this section is to show that each density $\rho\in\mathscr{X}_{>0}$ has only one subgradient. We do that by using a standard result from convex analysis and by the usual Hohenberg–Kohn argument.
@@
We now have all the necessary ingredients to show Gâteaux differentiability of the thermal universal functional.
@@
The thermal universal functional $F_{\mathrm{DM}}^{\beta}$ is Gâteaux differentiable at any density $\rho\in\mathscr{X}_{>0}$.
@@
According to Theorem 7, the functional $F_{\mathrm{DM}}^{\beta}$ is continuous on $\mathscr{X}_{>0}$ and $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is non-empty for all densities $\rho\in\mathscr{X}_{>0}$. By Lemma 13 it is then enough to show that $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is single-valued. Let us assume that there is a $\rho\in\mathscr{X}_{>0}$ for which $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is not single-valued. In particular, we assume that $v_{1},v_{2}\in-\partial F_{\mathrm{DM}}^{\beta}(\rho)$ with $v_{1}\neq v_{2}$. By Theorem 11 the two Gibbs states $\Gamma_{1}$ and $\Gamma_{2}$ corresponding to $v_{1}$ and $v_{2}$ both yield the density $\rho$. Moreover, Proposition 3 states $\Gamma_{1}\neq\Gamma_{2}$. We then argue like in the proof of the Hohenberg–Kohn theorem [21] that
$\Omega^{\beta}(v_{1})$ $=\Omega^{\beta}_{v_{1}}(\Gamma_{1})=\Omega^{\beta}_{v_{2}}(\Gamma_{1})+\langle\rho,v_{1}-v_{2}\rangle>\Omega^{\beta}_{v_{2}}(\Gamma_{2})+\langle\rho,v_{1}-v_{2}\rangle=\Omega^{\beta}(v_{2})+\langle\rho,v_{1}-v_{2}\rangle.$ (39)
Changing the roles of $v_{1}$ and $v_{2}$ and adding the two inequalities gives the contradiction
$\Omega^{\beta}(v_{1})+\Omega^{\beta}(v_{2})<\Omega^{\beta}(v_{2})+\Omega^{\beta}(v_{1}).$ (40)
Hence, $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ must be single-valued. ∎
@@ VIII Summary and outlook
Let $\rho\in\mathscr{X}_{>0}$, then Theorem 7 shows that $\partial F^{\beta}_{\mathrm{DM}}(\rho)$ is non-empty. This implies that $\rho$ is a minimizer for the variational principle in Eq. (22a), where the potential is in the negative subdifferential, i.e., $v\in-\partial F^{\beta}_{\mathrm{DM}}(\rho)$. The minimizer of Eq. (22b), guaranteed to exist by Theorem 11, is equal to the Gibbs state $\Gamma_{v}$. We conclude, that $\rho$ is $v$-representable. On the other hand, every density of a Gibbs state $\Gamma_{v}$ with $v\in H^{-1}(\mathbb{T})$ is contained in $\mathscr{I}$ according to Proposition 4 and is strictly positive by Theorem 12. In particular, every $v$-representable density is in $\mathscr{X}_{>0}$ and $F^{\beta}_{\mathrm{DM}}$ is not differentiable at densities in $\mathscr{I}\setminus\mathscr{X}_{>0}$. On the other hand, Theorem 14 shows that the thermal universal functional $F^{\beta}_{\mathrm{DM}}$ is Gâteaux differentiable at any $\rho\in\mathscr{X}_{>0}$. ∎
@@
In this work we provided an explicit characterization of the set of uniquely $v$-representable densities on the one-dimensional torus at elevated temperatures and established (Gâteaux) differentiability of the thermal universal functional. In the same manner the theory can be developed for any other interval with Neumann boundary conditions. As an extensive parameter, the volume needs to be finite, and thus a periodic setting is the natural choice for elevated temperatures. The restriction to one dimension remains as the main limitation of the current formulation even though some of the results like the uniqueness of the Gibbs state, strict positivity of the density and the uniqueness of the potential also hold in higher dimensions. Note that we cannot conclude Gâteaux differentiability in higher dimensions since continuity of the thermal universal functional might not be guaranteed. Another interesting observation is that we need to include distributional potentials for achieving a $v$-representability result already in one dimension. In DFT often the potential set $L^{3/2}+L^{\infty}$ is considered [28]. However, for zero temperature and single particles it was already observed by Englisch and Englisch [11] and Chayes et al. [10] that certain densities require distributional potentials for their realization. Sutter et al. [38] showed that distributional potentials of the form Eq. (3b) are sufficient for $v$-representability of the set $\mathscr{X}_{>0}$. For elevated temperature we further show that such distributional potentials are also necessary. Since it cannot be expected that in higher dimensions the space of necessary potentials gets more regular, we conjecture that distributional potentials (probably of a higher class) are also required there if a similar proof of $v$-representability is feasible. However, in the current construction, the continuous Sobolev embedding $H^{1}\hookrightarrow L^{\infty}$ is crucial and it holds in one dimension. For higher dimensions, we either need to require higher regularity or higher integrability conditions. This means that for densities we consider the Sobolev space $W^{m,p}$ and need to have $m$ and/or $p$ large enough. The problem is then that not all Gibbs state densities might be contained in $W^{m,p}$ and that a Hamiltonian with a potential in $W^{-m,q}=(W^{m,p})^{\ast}$ will usually not be self-adjoint [20]. Another open question is how to extend the results to the grand canonical ensemble of variable particle number. In this case, the particle-number operator is not bounded anymore and this is currently needed in the proof of Theorem 11, where we show that the weak-* limit yields the same density as the sequence. Additionally, we do not know if the square root of the Gibbs state density is in $H^{1}$. However, we think that the rest of the theory does not fundamentally change for the grand canonical ensemble and that we retrieve the same conclusions as for the canonical ensemble.
@@ Acknowledgements.
We are grateful to Thiago Carvalho Corso for helpful feedback on a previous version of this article. MP acknowledges support from the German Research Foundation under Grant SCHI 1476/1-1 and from ERC-2021-STG grant agreement No. 101041487 REGAL. The authors SMS and KJHG thank The Netherlands Organization for Scientific Research, NWO, for its financial support under Grant No. OCENW.KLEIN.434 and Vici Grant No. 724.017.001. RvL acknowledges the Academy of Finland grant under project number 356906. MR acknowledges the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG), EXC 2056, project ID 390715994.
@@ Appendix A Proof of Proposition 3
We still need to show that Proposition 3 holds. The result that the unique minimizer of the Helmholtz free energy is given by the Gibbs state $\Gamma_{v}=\mathrm{e}^{\beta H_{v}}/Z(v)$ and the relation $\Omega^{\beta}_{v}(\Gamma_{v})=-\beta\log Z(v)$ is shown in standard literature. We give here a more elaborated version than our previous works [17; 37] in the form of a formal proof without going into mathematical subtleties. A classical proof can be found in Mermin [30, Appendix] for the grand canonical ensemble but works analogously for the canonical one.
@@
To determine the minimizer of the Helmholtz free energy $\Omega^{\beta}_{v}(\Gamma)=\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$ we make it stationary with respect to small perturbations $\delta\Gamma$ that obey $\operatorname{Tr}\{\delta\Gamma\}=0$. We compute
@@
For the last remaining part of the proposition we need to show that $Z(v)$ is indeed finite for any potential $v\in H^{-1}(\mathbb{T})$. First, we show that the partition function is finite for a non-interacting system with zero potential, i.e., $H_{v}=T$. We denote this partition function by $Z_{0}(0)$. The eigenstates of the kinetic-energy operator $T=-\frac{1}{2}\Delta$ on $\mathbb{T}$ are Slater determinants built from the functions $\mathrm{e}^{\mathrm{i}2\pi p\cdot x}$ with $p\in\mathbb{Z}$ which have eigenvalues $2\pi^{2}p^{2}$. We estimate
@@
To show that $Z(v)$ is finite for a general interaction and a general potential we first show that $\Omega^{\beta}(v)$ is finite. For that recall that we assume that the interaction satisfies the KLMN condition from Eq. (10). Moreover, Sutter et al. [38, Lem. 16] showed that any potential $v\in H^{-1}$ also fulfills the KLMN condition $|V(\Psi,\Psi)|\leq a^{\prime}T(\Psi)+b^{\prime}$ where $a^{\prime}>0$ can be made arbitrary small. Here $V(\cdot,\cdot)$ is the quadratic form produced by $v\in H^{-1}$. Combing these two KLMN conditions guarantees the existence of new constants $0\leq a<1$ and $b\geq 0$ such that $\langle\Psi,H_{v}\Psi\rangle\geq(1-a)T(\Psi)-b$ for all $\Psi\in\mathscr{W}$. This condition translates to density matrices and we get
$\Omega^{\beta}(v)=\inf_{\Gamma\in\mathscr{D}}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log(\Gamma))\}\geq(1-a)\inf_{\Gamma\in\mathscr{D}}\operatorname{Tr}\{\Gamma(T+\big(\beta(1-a)\big)^{-1}\log(\Gamma))\}-b=-\beta^{-1}\log\big(Z_{0}(0)\big)-b>-\infty.$ (45)
Let $|H_{v}=\sum_{j=1}^{\infty}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of the Hamiltonian $H_{v}$ which exist due to the compact embedding of $Q(H_{v})=\big(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\big)^{\wedge N}$ into $\big(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2}\big)^{\wedge N}$. We define a sequence of density matrices by
@@
The left-hand side is bounded from below by Eq. (45) and we get that $Z_{n}[v]$ is uniformly bounded from above. This implies $Z(v)<\infty$.
Finally, to show that $\Gamma_{v}\in\mathscr{D}$ we need to show $S(\Gamma_{v})<\infty$. We use again $\log\Gamma_{v}=-\beta H_{v}-\log Z(v)$ and show with Lemma 16 that
$S(\Gamma_{v})=-\operatorname{Tr}\{\Gamma_{v}\log\Gamma_{v}\}=\beta\operatorname{Tr}\{\Gamma_{v}H_{v}\}+\log Z(v)\operatorname{Tr}\{\Gamma_{v}\}<\infty.$ (48)